# Computing a Fixed Point of Contraction Maps in Polynomial Queries

### v1

@@ 7 Promise Problem versus Total Search Version
[…]
@@
[…]
@@
Let $\left\{q^{1},\cdots,q^{m}\right\}$ be a set of points in $[0,1]^{k}$ and $\left\{a^{1},\cdots,a^{m}\right\}$ be the corresponding answers from the black-box oracle. There is a $(1-\gamma)$-contraction $f$ that is consistent with all the answers if and only if there is no pair $t_{1},t_{2}$ such that $\|a^{t_{1}}-a^{t_{2}}\|_{\infty}>(1-\gamma)\|q^{t_{1}}-q^{t_{2}}\|_{\infty}$.
@@
[…]
8 Conclusions
We gave an algorithm for finding an $\varepsilon$-fixed point of a contraction (or non-expansive) map $f:[0,1]^{k}\mapsto[0,1]^{k}$ under the $\ell_{\infty}$ norm in polynomial query complexity. Contraction maps under the $\ell_{\infty}$ norm are especially important because several longstanding open problems from various fields can be cast in this framework. The main open question is whether our algorithm can be implemented to run also with polynomial time complexity, or alternatively if there is another general-purpose (black-box) algorithm for contraction maps that finds an approximate fixed point in polynomial time. Resolving positively this question would have tremendous implications.
Another natural open question is whether similar polynomial query bounds can be obtained for contraction maps under the $\ell_{1}$-norm or norms $\ell_{p}$ with $p>2$. Although $\ell_{\infty}$ seems to arise more in applications, understanding better contractions under other norms would also be useful.
@@ Acknowledgements
We thank STOC anonymous reviewers for their feedbacks on improving the presentation.

### v2

@@ 7 Promise Problem versus Total Search Version
The problem $\textsc{Contraction}_{\infty}(\varepsilon,\gamma,k)$ is a promise problem, where we want to compute an $\varepsilon$-fixed point of a given function $f$ with promise that $f$ is a $(1-\gamma)$-contraction. For a promise problem, one can define its total search version by asking to find a desired solution as in the promise problem, or a short violation certificate indicating that the given function doesn’t satisfy the promise. The total search version of $\textsc{Contraction}_{\infty}(\varepsilon,\gamma,k)$, denoted $\textsc{T-Contraction}_{\infty}(\varepsilon,\gamma,k)$ is naturally defined as the following search problem.
Definition 4 (Total search version $\textsc{T-Contraction}_{\infty}(\varepsilon,\gamma,k)$).
@@
Our theorem in this section shows $\textsc{T-Contraction}_{\infty}(\varepsilon,\gamma,k)$ admits the same query bounds as $\textsc{Contraction}_{\infty}(\varepsilon,\gamma,k)$.
@@
Let $\left\{q^{1},\cdots,q^{m}\right\}$ be a set of points in $0,1]^{k}$ and $\left\{a^{1},\cdots,a^{m}\right\}$ be the corresponding answers from the black-box oracle. There is a $(1-\gamma)$-contraction $f$ that is consistent with all the answers if and only if there is no pair $t_{1},t_{2}$ such that $\|a^{t_{1}}-a^{t_{2}}\|_{\infty}>(1-\gamma)\|q^{t_{1}}-q^{t_{2}}\|_{\infty}$.
@@
It follows from Lemma 11 that we can use any algorithm that can solve the promise problem $\textsc{Contraction}_{\infty}(\varepsilon,\gamma,k)$ to solve the total search version $\textsc{T-Contraction}_{\infty}(\varepsilon,\gamma,k)$ within the same number of queries (without loss of generality, we can assume that the algorithm for the promise problem $\textsc{Contraction}_{\infty}(\varepsilon,\gamma,k)$ always queries the approximate fixed point it finds). To see it, consider the sequence of query points $q^{1},\ldots,q^{m}$ generated by the algorithm and the corresponding answers $a^{1},\ldots,a^{m}$ of the black-box oracle. If $\|q^{i}-a^{i}\|_{\infty}\leq\epsilon$ for some $i$, then the algorithm will return the $\epsilon$-approximate fixed point $q^{i}$ and terminate. If there is a pair of generated queries that violate the contraction property, then the algorithm can return the violating pair and terminate. Suppose that neither of these events happens; since all pairs of queries satisfy the contraction property, there is a $(1-\gamma)$-contraction map $f$ that is consistent with all the queries. Then the algorithm for the promise problem $\textsc{Contraction}_{\infty}(\varepsilon,\gamma,k)$ fails to find an $\epsilon$-approximate fixed point for the black-box oracle function $f$ within the prescribed time bound, a contradiction. Therefore, given any black-box oracle, the algorithm will find either an $\epsilon$-approximate fixed point or a violating pair within the same number of queries $m=O(k\log(1/\epsilon))$ as the promise problem.
8 Conclusions and Discussions
We gave an algorithm for finding an $\varepsilon$-fixed point of a contraction (or non-expansive) map $f:[0,1]^{k}\mapsto[0,1]^{k}$ under the $\ell_{\infty}$ norm in polynomial query complexity. Contraction maps under the $\ell_{\infty}$ norm are especially important because several longstanding open problems from various fields can be cast in this framework. The main open question is whether there is a general-purpose (black-box) algorithm for contraction maps that finds an approximate fixed point in polynomial time. Resolving positively this question would have tremendous implications.
Our results provide the necessary first step towards the goal above since, in the black-box model, any polynomial-time algorithm must imply polynomial-query complexity. Nevertheless, we currently see obstacles to making our algorithm time-efficient. Namely, it is not clear how to efficiently find (or even verify) an approximate balanced point, whose existence guarantee is based on non-constructive arguments.
On the other hand, it is possible that finding an approximate fixed point of contraction maps is computationally hard (e.g., CLS-hard or UEOPL-hard). However, our result restricts what approaches could work in this direction. At a high level, such reduction must work in a white-box manner. To convey the interpretation more concretely, take the canonical CLS-complete problem Continuous-Localopt as an example. The problem takes two continuous functions, $p$ and $g$, as inputs and asks for an approximate local optimum of $p$ with respect to $g$; see its formal definition in [FGHS23]. It could be the case where the reduction takes the polynomial-size arithmetic circuits that represent $p$ and $g$ as inputs and simulates their gates to generate a contraction map $f$ such that any approximate fixed point of $f$ encodes an approximate local optimum of $p$ with respect $g$. In contrast, it cannot be the case where the reduction treats the arithmetic circuits for $p$ and $g$ as evaluation oracles. Specifically, it cannot produce a contraction map $f$ where the evaluation of $f$ requires only a polynomial number of calls to the oracles for $p$ and $g$, since there are exponential query lower bounds for the Continuous-Localopt problem [HY20].
An open question regarding the query complexity of $\textsc{Contraction}_{\infty}(\varepsilon,\gamma,k)$ is whether the upper bound $O(k\log(1/\varepsilon))$ is tight. Currently, we are not aware of lower bounds that are stronger than the standard $\Omega(\log(1/\varepsilon))$, which holds even when $k=1$. In an earlier version of this article, we asked the question of whether similar query bounds can be obtained for contraction maps under the $\ell_{1}$-norm or norms $\ell_{p}$ with $p>2$. A very recent unpublished paper [HLSW25] answered this question affirmatively by showing $O(k^{2}\log(1/\varepsilon\gamma)))$ query upper bounds for all norms $\ell_{p}$ with $p\in[1,\infty)\cup\left\{\infty\right\}$.
@@ Acknowledgements
The authors are very grateful to the editor and anonymous reviewers for their helpful feedback on improving the presentation.